# Correcting coherent quantum errors by going with the flow

### v1

@@ I Introduction
In this paper, we find: (1) the average logical failure “rate” can increase as a function of the number of error correction rounds if coherent errors are allowed to build constructively; (2) if the $X$ and $Z$ noise are comparable, the distance of the code is $3$ or more, and one performs passive error correction through virtual Pauli frame updates, then the coherent noise will not build constructively across error correction rounds; (3) by also initializing the logical qubit in a random Pauli frame and using a code distance greater than $3$, the performance of the logical qubit with correlated noise will essentially be the same as the performance with uncorrelated Pauli noise having the same process fidelity apart from constructive/destructive interference within each round of error correction. These findings are in contrast to local noise model threshold theorems Terhal and Burkard (2005); Aliferis et al. (2006); Preskill (2013) that are quadratically penalized, and they are generally applicable to stabilizer codes without being limited to the asymptotics of large code sizes. With a “lazy” strategy of initializing in a random Pauli frame and performing passive error correction (and assuming comparable $X$ and $Z$ noise), the penalty for having correlated noise is essentially eliminated (with exception to Hamiltonian errors that may add constructively within error correction cycles).
The rest of the paper is as follows: In section II we describe the noise models we study. In section III we analyze the failure probability of logical idle in the code capacity model in the context of active error correction. In section IV we analyze the benefits of initializing in a random codespace and performing “passive” error correction using Pauli frame updates. In section V we discuss different time-correlated noise models and caveats when syndrome extraction is performed using gates that have correlated noise (i.e., a circuit model description). In section VI we present simulations that corroborate our findings. In section VII we present our conclusions.
@@ II Syndrome extraction models and noise Types
Note that we could combine any of the SEMs with any of the noise types. Sections III and IV are discussed in the context of the CCM. Section V sketches out a generalization of our arguments to the CM. We present simulation results for the CCM in VI.
@@ IV.1 Random initial codespace
To be more concrete, consider two uncorrectable Pauli-$X$ errors with the same syndrome in the expansion of the coherent $X$-noise operator $\exp(iH_{\mbox{noise}})=\ldots a_{1}E_{1}^{X}+a_{2}E_{2}^{X}\ldots$. Since $E_{1}^{X}$ and $E_{2}^{X}$ are uncorrectable and have the same syndrome it follows that $\pm E_{1}^{X}E_{2}^{X}$ is a stabilizer. If $\pm E_{1}^{X}E^{X}_{2}$ is a stabilizer then the erroneous post-correction state given this syndrome will have the term $(a_{1}\pm a_{2})\bar{X}|\bar{\psi}\rangle$ from these contributions. This error contribution can be impacted by the codespace to the extent that $|a+b|$ differs from |ab||a-b|. Initializing a logical qubit into a random codespace can be straightforward. For example, in a qLDPC code one may simply initialize physical qubits at random in $\{{|}0{\rangle},{|}1{\rangle}\}$ or $\{{|}+{\rangle},{|}-{\rangle}\}$, depending on ${|}\bar{0}{\rangle}$ or ${|}\bar{+}{\rangle}$ preparation, and start applying rounds of syndrome extraction to these physical qubits and the state will collapse into a random codespace. The syndrome measurements as well as the initial physical states indicate which one. Our simulations on the rotated surface code in Sec. VI show that there is a substantial benefit from starting in a random codespace for combating Hamiltonian noise.
@@ V.2 Circuit model (CM)
VI Simulation Results
@@
In Fig. (2) we simulate the medial surface code in the CCM with different combinations of global continuous noise with a constant $\epsilon$ (equivalent to local noise with $\sigma_{\epsilon}=0$ and $\mu_{\epsilon}=\epsilon$) and discrete Pauli flips of probability $p$. Subscripts of $x$ and $z$ in the legend denote $X$-type and $Z$-type noise. The plotted failure probabilities only account for the effects from the $X$ noise. Where present, the discrete Pauli-$Z$ flips are tracked and perfectly corrected at the end so that we can understand its effect purely on the continuous $X$ noise, testing the random walk described in Sec. IV.2 that dampens the coherent build-up of error probability due to the continuous noise. These are introduced at probability $p=\epsilon^{2}$ which is the same probability as an individual spin flip after one application of the Hamiltonian noise with strength $\epsilon$. That is, these continuous and discrete noise contributions are equivalent apart from coherent noise effects introduced in the continuous noise component. Note that the $X$ and $Z$ labels are symmetric for this consideration (the results would be the same if we had continuous $Z$ noise and discrete $X$ noise). The simulations with nonzero $p_{z}$ in Fig. (2) mimic the case of having both $X$ and $Z$ type continuous noise of comparable strength but isolate the effects of each one indepenently.
@@
The worst case scenario in Fig. (2) is when the logical qubit is initialized in the syndrome-zero codespace and there are no Pauli-$Z$ flips ($p_{z}=0$). This case exhibits a prominent quadratic scaling of the failure probability as a function of the number of cycles $n$. It has the worst case initial codespace (the syndrome-zero subspace) and has no $Z$ flips to induce a random walk effect. Introducing Pauli-$Z$ flips that are actively corrected is an equivalent model (by correcting flips, you lose the benefits). Introducing tracked, discrete Pauli-$Z$ flips that are corrected only at the end, we observe benefits of a random walk effect. Initializing into the syndrome-zero codespace with $p_{z}=\epsilon_{x}^{2}$, we see the same initial steep slope of $P_{f}$ vs $n$ but a downward curvature due to the benefits of a random walk in the codespace. There is better initial performance when we start in a random codespace. However, without introducing Pauli-$Z$ flips, we see a quadratic upward curve in $P_{f}$ vs $n$. Finally, when we use a random initial codespace and have the Pauli-$Z$ flips, we see the best performance without much upward curvature. For comparison, we also show the case when the $X$ noise is discrete. The difference in these final best cases shows a risidual loss of performance due to continuous noise that is expected at $d=3$ according to Eq. (32) with $\sigma_{\epsilon}=0$. At higher distances, the continuous noise penalty will be suppressed even further. This comparison is also an indication that the perturbative approximation is appropriate even at a relatively high value of $\epsilon=0.1$.

### v2

# arXiv:2602.21076v2[quant-ph] 02 Mar 2026

@@ I Introduction
In this paper, we find: (1) the average logical failure “rate” can increase as a function of the number of error correction rounds if coherent errors are allowed to build constructively; (2) if the $X$ and $Z$ noise are comparable and one performs passive error correction through virtual Pauli frame updates, then the coherent noise will not build constructively across error correction rounds; (3) by also initializing the logical qubit in a random Pauli frame, the performance of the logical qubit with correlated noise will essentially be the same as the performance with uncorrelated Pauli noise having the same process fidelity apart from constructive/destructive interference within each round of error correction. These findings are in contrast to local noise model threshold theorems Terhal and Burkard (2005); Aliferis et al. (2006); Preskill (2013) that are quadratically penalized, and they are generally applicable to stabilizer codes without being limited to the asymptotics of large code sizes. With a “lazy” strategy of initializing in a random Pauli frame and performing passive error correction (and assuming comparable $X$ and $Z$ noise), the penalty for having correlated noise is essentially eliminated (with exception to Hamiltonian errors that may add constructively within error correction cycles).
The rest of the paper is as follows: In section II we describe the noise models we study. In section III we analyze the failure probability of logical idle in the code capacity model in the context of active error correction. In section IV we analyze the benefits of initializing in a random codespace and performing “passive” error correction using Pauli frame updates. In section V we discuss different time-correlated noise models and caveats when syndrome extraction is performed using gates that have correlated noise (i.e., a circuit model description). In section VI we present simulations, corroborating our findings, for code capacity models.In section VII we present our conclusions.
@@ II Syndrome extraction models and noise Types
Note that we could combine any of the SEMs with any of the noise types. Sections III and IV are discussed in the context of the CCM. Section V sketches out a generalization of our arguments to the CM. We present simulation results for the CCM in Sec. VI.
@@ IV.1 Random initial codespace
To be more concrete, consider two uncorrectable Pauli-$X$ errors with the same syndrome in the expansion of the coherent $X$-noise operator $\exp(iH_{\mbox{noise}})=\ldots a_{1}E_{1}^{X}+a_{2}E_{2}^{X}\ldots$. If $E_{1}^{X}E^{X}_{2}$ is a stabilizer then the erroneous post-correction state given this syndrome will have the term $(a_{1}+a_{2})\bar{X}|\bar{\psi}\rangle$ from these contributions. This error contribution can be impacted by the codespace to the extent that $|a_{1}+a_{2}|$ differs from $|a_{1}-a_{2}|$. Initializing a logical qubit into a random codespace can be straightforward. For example, in a qLDPC code one may simply initialize physical qubits at random in $\{{|}0{\rangle},{|}1{\rangle}\}$ or $\{{|}+{\rangle},{|}-{\rangle}\}$, depending on ${|}\bar{0}{\rangle}$ or ${|}\bar{+}{\rangle}$ preparation, and start applying rounds of syndrome extraction to these physical qubits and the state will collapse into a random codespace. The syndrome measurements as well as the initial physical states indicate which one. Our simulations on the rotated surface code in Sec. VI show that there is a substantial benefit from starting in a random codespace for combating Hamiltonian noise.
@@ V.2 Circuit model (CM)
VI Code capacity model Results
@@
In Fig. (2) we simulate the medial surface code in the CCM with different combinations of global continuous noise with a constant $\epsilon$ (equivalent to local noise with $\sigma_{\epsilon}=0$ and $\mu_{\epsilon}=\epsilon$) and discrete Pauli flips of probability $p$. Subscripts of $x$ and $z$ in the legend denote $X$-type and $Z$-type noise. The plotted failure probabilities only account for the effects from the $X$ noise. Where present, the discrete Pauli-$Z$ flips are tracked and perfectly corrected at the end so that we can understand its effect purely on the Hamiltonian $X$ noise, testing the random walk described in Sec. IV.2 that dampens the coherent build-up of error probability due to the Hamiltonian noise. These are introduced at probability $p=\epsilon^{2}$ which is the same probability as an individual spin flip after one application of the Hamiltonian noise with strength $\epsilon$. That is, these continuous and discrete noise contributions are equivalent apart from coherent noise effects introduced in the Hamiltonian noise component. Note that the $X$ and $Z$ labels are symmetric for this consideration (the results would be the same if we had continuous $Z$ noise and discrete $X$ noise). The simulations with nonzero $p_{z}$ in Fig. (2) mimic the case of having both $X$ and $Z$ type Hamiltonian noise of comparable strength but isolate the effects of each one independently.
@@
The worst case scenario in Fig. (2) is when the logical qubit is initialized in the syndrome-zero code space and there are no Pauli-$Z$ flips ($p_{z}=0$). This case exhibits a prominent quadratic scaling of the failure probability as a function of the number of cycles $n$. It has the worst case initial code space (the syndrome-zero subspace) and has no $Z$ flips to induce a random walk effect. Introducing Pauli-$Z$ flips that are actively corrected is an equivalent model (by correcting flips, you lose the benefits). Introducing tracked, discrete Pauli-$Z$ flips that are corrected only at the end, we observe benefits of a random walk effect. Initializing into the syndrome-zero code space with $p_{z}=\epsilon_{x}^{2}$, we see the same initial steep slope of $P_{f}$ vs $n$ but a downward curvature due to the benefits of a random walk in the codespace. There is better initial performance when we start in a random codespace. However, without introducing Pauli-$Z$ flips, we see a quadratic upward curve in $P_{f}$ vs $n$. Finally, when we use a random initial codespace and have the Pauli-$Z$ flips, we see the best performance without much upward curvature. For comparison, we also show the case when the $X$ noise is discrete. The difference in these final best cases shows a residual loss of performance due to continuous noise that is expected at $d=3$ according to Eq. (32) with $\sigma_{\epsilon}=0$. At higher distances, the continuous noise penalty will be suppressed even further. This comparison is also an indication that the perturbative approximation is appropriate even at a relatively high value of $\epsilon=0.1$.